# miniCTX: Neural Theorem Proving with

### v1

@@ 6 Related Work
Formal theorem proving with language models.
GPT-f [9] pioneered the use of language models for theorem proving via next tactic prediction given the current proof state, a technique adopted by many subsequent methods [24, 7, 19, 22, 1, 23]. ReProver [8] conditions each generation on retrieved premises, while Draft-sketch-prove [25] conditions each generation on an informal proof. Baldur [26] fine-tunes a model with 50 lines of the preceding file content as context, but unlike file-tuning trains the model to generate a full proof without proof states. More broadly, machine learning for formal theorem proving is an active research area; see [27, 28] for surveys.
Theorem proving data extraction. Several tools extract training data from interactive theorem provers, including CoqGym [29] for Coq, PISA [24] for Isabelle, LeanStep [7] for Lean 3, and LeanDojo [8] for Lean 3 and 4. Recently, lean-training-data [20] provides tools for extracting proof states and other information using Lean 4 metaprogramming, which we anecdotally found to be easiest to modify and fastest among Lean 4 data extraction tools. Our ntp-toolkit adds 3 new tools on top of this code, along with a pipeline for running on any Lean projects and instruction tuning.
Theorem proving benchmarks. Theorem proving methods are typically evaluated in two settings: (1) standalone competition problems [6] or textbook [30] problems; (2) holding out theorems from a mathematical library that the model is trained on, such as Mathlib for Lean  [7, 22, 8] or the Archive of Formal Proofs for Isabelle [24, 26]. The first does not test the use of context, while the second tests only theorem-level generalization. miniCTX is designed to test the use of context as well as theorem-level, context-level, and project-level generalization across several mathematical domains.
@@ A.1 Example Entry
"file": "(filename)",
@@
# Whether the theorem uses definitions or lemmas defined in this file
@@
# Whether the theorem uses definitions or lemmas defined in another file in this repository
"repositoryPremises": false
@@
A.2 PrimeNumberTheoremAnd Example
@@ B.1 Math2001
Math2001  [15] contains the Lean code for the book The Mechanics of Proof by Heather Macbeth, an introductory text on mathematical theorem proving with accompanying Lean code. Each chapter of The Mechanics of Proof covers an introductory topic and walks through how to write the associated mathematics in Lean, along with exercises. The topics include proofs by calculation, proofs with structure, parity and divisibility, logic, induction, number theory, functions, sets, and relations. A unique aspect of Math2001 is that it disables common Lean automation for pedagogical purposes. For example, a student must write out an equality proof in detail, with each step justified. It also defines new tactics and definitions separate from the common Lean libraries. Typically a file in the textbook will show examples of such proofs, followed by exercises for a student to complete. We can view this as a form of contextual adaptation: a model must prove the theorem according to the constraints of the textbook. Math2001 has 41 files that include examples and exercises. We selected 1 to 2 theorems from each file (depending on the length of the file), for a total of 50 theorems. Of these, 31 have no proof in the Math2001 repository, hence testing theorem-level generalization.
@@ C.1 Data extraction
ntp-toolkit contains a general-purpose data extraction tool that extracts examples from an arbitrary Lean 4 repository and formats them into examples for language-model fine-tuning. The tool is implemented in Lean based on Kim Morrison’s lean-training-data.
@@ Appendix D Additional results and analysis
D.1 Label distribution
Figure 4: Percentage of different dependencies in miniCTX
D.2 Performance by proof length and context length.
Figure 5: Performance of two methods on different difficulty levels and context lengths
D.3 Example of modifying proofs from context
@@
D.4 Example of using theorems from context
@@
D.5 Example of using unseen definitions
The State-tactic-tuned model is able to utilize the even unseen definitions that appear in the proof state. In the following example Set.uIoo is a newly defined definition, which is never seen for state-tactic tuned model:
@@ D.5 Example of using unseen definitions
Appendix E Dataset Hosting, Licensing, and Maintenance
miniCTX is released on HuggingFace: l3lab/miniCTX [^5] Croissant metadata record: https://huggingface.co/api/datasets/l3lab/miniCTX/croissant., distributed under the Apache 2.0 license. Data extraction tool ntp-toolkit is released on GitHub: cmu-l3/ntp-toolkit, under the MIT license. We note that the underlying data for the individual splits of miniCTX are also released under the Apache 2.0 license. We include the licensing information in the dataset repository. We plan to regularly update and maintain the dataset to include examples from new projects.
Appendix F Annotation Guideline
With the NTP-TOOLKIT, users can extract and annotate new theorems from any valid Lean project.
@@ F.1 Preliminary
The evaluation code relies heavily on the Lean REPL, which operates within the project environment. Therefore, it is essential that the project builds without any errors. Additionally, the version of Lean used in the project should match the version supported by the REPL. While the Lean REPL supports versions $\geq$ 4.3.0, for the best experience, we recommend evaluating projects that use Lean versions between 4.5.0 and 4.7.0.
@@ F.2 Using the NTP-TOOLKIT
The NTP-TOOLKIT is designed to facilitate the extraction and annotation of theorem proving data from Lean projects. To use the toolkit for annotating theorems, follow these steps:
@@
    Installation: Clone the NTP-TOOLKIT repository from GitHub to your local machine. Ensure that you have the required dependencies installed, as listed in the repository’s README file.
@@
    Configuration: Configure the toolkit to point to the directory of your Lean project. Make sure that your project is using a compatible version of Lean as per the guidelines in the previous section.
@@
    Data Extraction: Run the data extraction script provided by the toolkit. Specify the full_proof_training_data option to focus the extraction on datasets suitable for training models with full proofs.

### v2

@@ 6 Related work
Formal theorem proving with language models. GPT-f (Polu & Sutskever, 2020) pioneered the use of language models for theorem proving via next tactic prediction given the current proof state, a technique adopted by many subsequent methods (Jiang et al., 2021; Han et al., 2022; Lample et al., 2022; Polu et al., 2022; Welleck & Saha, 2023; Azerbayev et al., 2024). ReProver (Yang et al., 2023) conditions each generation on retrieved premises, while Draft-sketch-prove (Jiang et al., 2023) conditions each generation on an informal proof. Baldur (First et al., 2023) fine-tunes a model with 50 lines of the preceding file content as context, but unlike file-tuning trains the model to generate a full proof without proof states. More broadly, machine learning for formal theorem proving is an active research area; see Lu et al. (2023); Li et al. (2024) for surveys.
Theorem proving data extraction. Several tools extract training data from interactive theorem provers, including CoqGym (Yang & Deng, 2019) for Coq, PISA (Jiang et al., 2021) for Isabelle, LeanStep (Han et al., 2022) for Lean 3, and LeanDojo (Yang et al., 2023) for Lean 3 and 4. Recently, lean-training-data (Morrison, 2023) provides tools for extracting proof states and other information using Lean 4 metaprogramming, which we anecdotally found to be easiest to modify and fastest among Lean 4 data extraction tools. Our ntp-toolkit adds 3 new tools on top of this code, along with a pipeline for running on any Lean projects and instruction tuning.
Theorem proving benchmarks. Theorem proving methods are typically evaluated in two settings: (1) standalone competition problems (Zheng et al., 2022) or textbook (Azerbayev et al., 2023) problems; (2) holding out theorems from a mathematical library that the model is trained on, such as Mathlib for Lean  (Han et al., 2022; Polu et al., 2022; Yang et al., 2023) or the Archive of Formal Proofs for Isabelle (Jiang et al., 2021; First et al., 2023). The first does not test the use of context, while the second tests only theorem-level generalization. miniCTX is designed to test the use of context as well as theorem-level, context-level, and project-level generalization across several mathematical domains.
Premise selection. Premise selection is an extensively studied class of methods for handling context in theorem proving. These methods retrieve useful lemmas from previously proved results, to help prove the current theorem. Many recent premise retrievers embed theorems and premises to a common space, and then taking a similarity measure (Yang et al., 2023; Mikuła et al., 2023), or use a classifier to determine if a premise is relevant (Irving et al., 2016; Han et al., 2022). Graph2Tac (Blaauwbroek et al., 2024) proposes an online learning method that can learn from both proofs and premises in a new context, which is particularly relevant to our work since it tackles project-level generalization. However, such methods cannot combine with more advanced pre-training-based methods without risking data contamination during evaluation. miniCTX aims to fill this gap by our temporal split.
@@ 7 Conclusion
Acknowledgements
Sean Welleck thanks Convergent Research, the Lean FRO, and the OpenAI Researcher Access Program for their support.
@@ A.1 Example Entry
# Source metadata
"file": "MyProject/Square.lean",
"module": "MyProject.Square",
@@
# Number of definitions or lemmas defined in this file that the theorem uses
@@
"numInFilePremises": 1,
# Number of definitions or lemmas defined in this repository that the theorem uses (including in-file ones)
"repositoryPremises": true
"numRepositoryPremises": 1,
# Number of total premises (in file, repository, or otherwise)
"numPremises": 2,
# Modules imported in the current file
"importedModules": ["Mathlib.Data.Real.Basic", ...]
# Proof metadata
@@
A.2 Prime Number Theorem example
@@ B.1 Math2001
Math2001 (Macbeth, 2023) contains the Lean code for the book The Mechanics of Proof by Heather Macbeth, an introductory text on mathematical theorem proving with accompanying Lean code. Each chapter of The Mechanics of Proof covers an introductory topic and walks through how to write the associated mathematics in Lean, along with exercises. The topics include proofs by calculation, proofs with structure, parity and divisibility, logic, induction, number theory, functions, sets, and relations. A unique aspect of Math2001 is that it disables common Lean automation for pedagogical purposes. For example, a student must write out an equality proof in detail, with each step justified. It also defines new tactics and definitions separate from the common Lean libraries. Typically a file in the textbook will show examples of such proofs, followed by exercises for a student to complete. We can view this as a form of contextual adaptation: a model must prove the theorem according to the constraints of the textbook. Math2001 has 41 files that include examples and exercises. We selected 1 to 2 theorems from each file (depending on the length of the file), for a total of 50 theorems. Of these, 31 have no proof in the Math2001 repository, hence testing theorem-level generalization.
@@ C.1 Data extraction
ntp-toolkit contains a general-purpose data extraction tool that extracts examples from an arbitrary Lean 4 repository and formats them into examples that can be used to compile miniCTX, as well as for language-model fine-tuning. The tool is implemented in Lean based on Kim Morrison’s lean-training-data.
@@ Appendix D Additional results and analysis
D.1 Dependency distribution
Figure 5: Percentage of different dependencies in the human-written proof of theorems in miniCTX
D.2 Performance by proof length and context length
Figure 6: Performance by dependency type. For each theorem in miniCTX, we record as metadata whether its human-written proof depends on other definitions or theorems in the same file, and test the performance of baselines on each type. File-tuned models substantially outperform state-tactic tuned models on theorems with definition and/or theorem dependencies.
Figure 7: Performance of two baselines on different difficulty levels and context lengths, as measured by the length of human-written proof in lines and the size of the preceding file contents in tokens. File-tuning substantially improves theorem-proving abilities across all cases, but especially when the theorem is easier and the context is longer.
D.3 Interference between in-file context and retrieved premises
In our experiments, we attempted to supply both in-file context (in the form of preceding code) and premise context (in the form of retrieved premises) to GPT-4o for proving a theorem. In Figure 8, we present an analysis of the impact of the length of retrieved premises on the resulting proof success rate.
Longer retrieved premises hurt performance.
The results indicate that problems with a lower premise-to-context length ratio tend to have higher success rates. Specifically, successful problems often feature relatively shorter premises as proportion of the full context length. This suggests that models are better able to utilize and focus on relevant in-file context when the cross-file premises are proportionally smaller. Conversely, when the length of the premises becomes relatively large compared to the full context, it may overwhelm or distract the model, reducing its ability to effectively utilize the in-file information. This finding highlights the importance of ensuring a balanced integration of premises with the in-file context to maintain model focus and improve proof generation performance.
Figure 8: Impact of length of retrieved premises on GPT-4o model performance. A higher premise-to-context length ratio is correlated with lower success rates, suggesting that too much premise in context overwhelms the model.
D.4 Example of learning proofs from context
@@
D.5 Example of using theorems from context
@@
D.6 Example of using unseen definitions
The state-tactic tuned model is able to utilize the unseen definitions that appear in the proof state. In the following example Set.uIoo is a newly defined definition, which is never seen for state-tactic tuned model:
@@ D.6 Example of using unseen definitions
This leads the state-tactic tuned model to perform similarly well as the file-tuned model on problems that only rely on new definitions, and not new theorems (see Figure 6).
Appendix E Dataset hosting and maintenance
miniCTX is released on HuggingFace: l3lab/miniCTX, distributed under the Apache 2.0 license. Data extraction tool ntp-toolkit is released on GitHub: cmu-l3/ntp-toolkit, under the MIT license. We note that the underlying data for the individual splits of miniCTX are also released under the Apache 2.0 license. We include the licensing information in the dataset repository. We plan to regularly update and maintain the dataset to include examples from new projects.
Appendix F NTP-Toolkit guideline
We introduced ntp-toolkit in §3.4. With the ntp-toolkit, users can extract and annotate new theorems and proofs from any valid Lean project, in miniCTX format. The extracted data can be used either as updates to miniCTX, or as training data (for which we also provide instruction tuning utilities). We also develop a lightweight evaluation framework for easy evaluation on miniCTX.
@@ F.1 Preliminary
The evaluation code relies heavily on the Lean REPL (Lean Prover Community, 2024), which operates within the project environment. Therefore, it is essential that the project builds without any errors. Additionally, the version of Lean used in the project should match the version supported by the REPL. While the Lean REPL supports versions $\geq$ 4.3.0, for the best experience with data extraction and evaluation, we recommend evaluating projects that use Lean version 4.7.0 (all miniCTX theorems are in 4.7.0). We plan to continuously update ntp-toolkit to support newer versions.
@@ F.2 Using the ntp-toolkit
The ntp-toolkit is designed to easily extract and annotate theorem proving data from Lean projects, by simply providing the project URL. To use the ntp-toolkit for data extraction, follow these steps:
@@
    Installation: Clone the ntp-toolkit repository from GitHub to your local machine, and checkout the Lean version tag corresponding to the extracted project (e.g., v4.7.0). Ensure that you have the required dependencies installed, as listed in the repository’s README file.
@@
    Configuration: Supply GitHub URL, commit hash, and root modules of your Lean project in a JSON configuration file. Make sure that your project is using a compatible version of Lean. ntp-toolkit will extract data from all modules imported by the root modules.
@@
    Data extraction: Run the data extraction script provided by the toolkit. Specify the --full_proof_training_data and --premises options to extract miniCTX-style data, which will be stored in an minictx.jsonl output file. Specify the --declarations option to additionally extract the premises in each module, for premise retrieval. The full_proof_training_data outputs can be additionally used for fine tuning (assuming the extracted data is dated before the current temporal split of miniCTX).